The demand curve for a public park for two consumers who represent society is given by:



๐‘ƒ = 150 โˆ’ ๐‘„๐ท1, ๐‘ƒ = 250 โˆ’ ๐‘„๐ท2



Graph the two demand curves and show the marginal social benefit curve for this public



park. If the marginal cost of providing the park was โ‚ฌ240, what would the optimum



provision of this park be? Explain why any quantity above or below this amount would



represent a less than efficient allocation.


1
Expert's answer
2021-11-26T12:23:46-0500

If ๐‘ƒ = 150 โˆ’ ๐‘„๐ท1, ๐‘ƒ = 250 โˆ’ ๐‘„๐ท2, then Qd1 = 150 - P, Qd2 = 250 - P.

Total demand is Qd = Qd1 + Qd2 = 400 - 2P.

MSB = Qd2 = 250 - P or P = 250 - Q.

If the marginal cost of providing the park was โ‚ฌ240, then:

MSB = MC,

250 - Q = 240,

Q = 10 units.